OpenStudy (agentnao):

What is the sum of a 12–term arithmetic sequence where the last term is 13 and the common difference is –10?

OpenStudy (angela210793):

a12=a1+(n-1)*d 13=a1+11*(-10) 13=a1-110 a1=123 Sn=(a1+a12)*12/2=(123+13)*6=816

OpenStudy (agentnao):

Thank you! You really are a superhero :)

OpenStudy (angela210793):

lol..no I'm not :) and yw :D

OpenStudy (akshay_budhkar):

angela your level tells it.. YOU ARE A SUPERHERO!